A diamond lies completely in a pool of water. A beam of light hits the surface of the diamond at an angle of 75 °. Determine the width of the angle at which the light beam continues in the diamond.
1
Expert's answer
2020-05-03T17:21:17-0400
"\frac{\sin{75}}{\sin{x}}=\frac{n_d}{n_w}=\frac{2.42}{1.33}"

"x=32\degree"